Calculate the pOH of a 1.21 × 10−3M HClO3 solution.

Report your answer to TWO places past the decimal.
Answer #1

Answer

pH = 2.92

Explanation

Chloric acid is a strong acid , dissociates completely

HClO3 -------> H+ + ClO3-

Concentration of HClO3 = 1.21×10-3M

Concentration of H+ = 1.21×10-3M

pH = - log[H+]

pH = - log(1.21×10-3)

pH = 2.92